# The Nonlinear Filter Model of Stream Cipher Redivivus

### v3

@@ Abstract
The nonlinear filter model is an old and well understood approach to the design of secure stream ciphers. Extensive research over several decades has shown how to attack stream ciphers based on this model and has identified the security properties required of the Boolean function used as the filtering function to resist such attacks. This led to the problem of constructing Boolean functions which provide adequate security and at the same time are efficient to implement. Unfortunately, over the last two decades no good solutions to this problem appeared in the literature. The lack of good solutions has effectively led to nonlinear filter model becoming more or less obsolete. This is a big loss to the cryptographic design toolkit, since the great advantage of the nonlinear filter model lies, beyond its simplicity and its ability to provide low-cost solutions for hardware-oriented stream ciphers, in the accumulated knowledge about the security requirements for tap positions and the filter function, which gives great confidence in its security when all criteria are met. In this paper we construct balanced functions on an odd number $n\geq 5$ of variables with the following provable properties: linear bias equal to $2^{-\lfloor n/2\rfloor-1}$, algebraic degree equal to $2^{\lfloor\log_{2}\lfloor n/2\rfloor\rfloor}$, algebraic immunity at least $\lceil(n-1)/4\rceil$, fast algebraic immunity at least $1+\lceil(n-1)/4\rceil$, and the functions can be implemented using $O(n)$ NAND gates. The functions are obtained from a simple modification of the well known class of Maiorana-McFarland bent functions. Due to the efficiency of implementation, for any target security level, we can construct efficient implementable functions which provide the required level of resistance to fast algebraic and fast correlation attacks. Previously known efficiently implementable functions have too large linear bias making them inappropriate even with a large number of variables. By appropriately choosing $n$ and the length $L$ of the linear feedback shift register, we show that it is possible to obtain examples of stream ciphers which are provably $\kappa$-bit secure against well known classes of attacks for various values of $\kappa$. We provide concrete proposals for $\kappa=80,128,160,192,224$ and $256$ using LFSRs of lengths 163, 257, 331, 389, 449, 521 and filtering functions on 75, 119, 143, 175, 203 and 231 variables. For the $80$-bit, $128$-bit, and the $256$-bit security levels, the circuits for the corresponding stream ciphers require about 1743.5, 2771.5, and 5607.5 NAND gates respectively. For the $80$-bit and the $128$-bit security levels, the gate count estimates compare quite well to the famous ciphers Trivium and Grain-128a respectively, while for the $256$-bit security level, we do not know of any other stream cipher design which has such a low gate count.
@@ 1 Introduction
The nonlinear filter model of stream ciphers is several decades old; one may note that the model was extensively discussed in the book by Rueppel [48] which was published in the mid-1980s. The nonlinear filter model consists of two components, namely a linear feedback shift register (LFSR) and a Boolean function $f$ which is applied to a subset of the bits of the LFSR at fixed positions (called tap positions). At each clock, $f$ is applied to the present state of the LFSR to produce a single keystream bit and simultaneously the LFSR also moves to the next state. LFSRs are very efficient to implement, particularly in hardware. So the implementation efficiency of the nonlinear filter model is essentially determined by the efficiency of implementation of $f$.
@@
A class of guess-then-determine attacks are known against nonlinear filter model. The inversion attack [27] is the first such attack. Subsequently, a number of papers have developed the idea into the generalised inversion attack, the filter state guessing attack, and the generalised filter state guessing attack [27, 28, 33, 46, 58]. Recommendations for resistance to the state guessing attack do not specify conditions on the filtering function. Rather, the recommendations specify conditions on the tap positions, i.e. the positions of the LFSR that are fed as input to the filtering function. We note that such recommendations implicity assume that the number of variables of the filtering function is much smaller than the desired security level.
@@
The well known Maiorana-McFarland class of bent functions is defined as follows. For $m\geq 1$, let $\mathbf{X}$ and $\mathbf{Y}$ be two vectors of $m$ variables. Then a $2m$-variable Maiorana-McFarland bent function is defined to map $(\mathbf{X},\mathbf{Y})$ to $\langle\pi(\mathbf{X}),\mathbf{Y}\rangle\oplus h(\mathbf{X})$, where $\pi$ is a bijection from $m$-bit strings to $m$-bit strings and $h$ is any $m$-variable Boolean function. It is well-known that the nonlinearity of Maiorana-McFarland functions does not depend on the choice of $h$ nor on that of the permutation $\pi$. Bad choices of both $\pi$ and $h$ (for example, $\pi$ to be the identity permutation and $h$ to be a constant function) provide functions whose algebraic immunity is low. On the bright side, we observe that it may be possible to improve the algebraic resistance by properly choosing $h$. Assume that $\pi$ is chosen to be the identity permutation. It is known [24] that the majority function possesses the best possible algebraic immunity. Motivated by this fact, we choose $h$ to be the majority function on $m$ variables. We prove that the resulting bent function on $2m$ variables has algebraic immunity at least $\lceil m/2\rceil$, and hence fast algebraic immunity at least $1+\lceil m/2\rceil$. On the implementation aspect, we show that the majority function can be computed using $O(m)$ NAND gates. So the obtained bent function has maximum nonlinearity, sufficient (fast) algebraic immunity when the number of variables is large enough (to protect against fast algebraic attacks at a specified security level), and at the same time is quite efficient to implement. Indeed, the (fast) algebraic immunity is not the maximum possible, but due to the implementation efficiency, it is possible to increase the number of variables to achieve the desired level of algebraic resistance. A novelty of our work is the observation and the proof that choosing $h$ to be the majority function improves the algebraic immunity. While Maiorana-McFarland bent functions have been extensively studied in the literature, this simple observation has escaped the notice of earlier researchers.
@@
One problem is that a bent function is not balanced. This problem is easily rectified by XORing a new variable to the bent function, obtaining a function on an odd number of variables. This modification requires only one extra XOR gate for implementation. In terms of security, the modification does not change the linear bias. We prove that the algebraic immunity of the modified function is at least the algebraic immunity of the bent function, and hence the fast algebraic immunity is at least one plus the algebraic immunity of the bent function. So the algebraic resistance of the modified function is essentially the same as that of the bent function. A consequence of unbalancedness of the filtering function is that the keystream sequence is unbalanced, which leads to a distinguishing attack. The simple modification of XORing an extra variable prevents such a distinguishing attack. The resistance to fast correlation attacks and fast algebraic attacks are ensured by the bent function (on an appropriate number of variables). A positive aspect of XORing a new variable is that it prevents (see Theorem 2 of [27]) certain kinds of information leakage which was identified in [3]. On the negative side, filter functions of the form $W+g(\mathbf{Z})$ were shown to be susceptible to inversion attack [27]. We prevent this attack by choosing the gap between the tap positions to be sufficiently large; this countermeasure was already proposed in [27]. More generally, we show that due to the fact that we use filter functions on a large number of variables, the various kinds of state guessing attacks [28, 33, 46, 58] do not apply to our proposals.
@@
We note that there are some old (prior to the advent of algebraic attacks) works [50, 51, 30] on efficient implementation of Boolean functions on a large number of variables targeted towards the nonlinear combiner model of stream ciphers. There are also a few later works [45, 18] on implementation of Maiorana-McFarland type functions. These works, however, do not cover the implementation of the functions that we consider, the reason being that these functions themselves do not appear earlier in the literature.
@@
Well known stream ciphers such as Trivium [9], Mickey [4], Grain-128a [1], SNOW [26] and Sosemanuk [5], use novel and ingenious ideas. Nonetheless, these are standalone designs. The nonlinear filter model, on the other hand, is a model for stream cipher design. Due to the simplicity of the nonlinear filter model, provable properties of the filtering function can be translated into provable resistance of the stream cipher against well known classes of attacks. In particular, for the proposals that we put forward, the provable linear bias of the filtering function translates to provable protection against a large class of fast correlation attacks, and the provable lower bound on the (fast) algebraic immunity translates to provable resistance against algebraic attacks. Stream ciphers based on either nonlinear finite state machines, or nonlinear feedback shift registers do not enjoy this advantage, i.e. for such stream ciphers it is very hard to obtain provable guarantees against various well known types of attacks. As an example, our proposal at the 128-bit security level generates keystream for which the best linear approximation has provable correlation of $2^{-60}$, while at the same security level, for Sosemanuk [5] the best known [39] linear approximation has correlation $2^{-20.84}$, and it is not known whether there are approximations with higher correlations. Similarly, the time complexities of fast algebraic attacks against Trivium, Mickey, Grain-128a, SNOW and Sosemanuk are not known. It is believed that these stream ciphers can withstand fast algebraic attacks. In contrast, we are able to prove that at the appropriate security levels, the fast algebraic attack is ineffective against the new nonlinear filter model based stream cipher proposals that we put forward.
One advantage of using LFSRs is that it is possible to provably ensure that the LFSR has a maximum period. For stream ciphers based on NFSRs, such as Trivium and Mickey, such provable assurance is not available. For stream ciphers which use a combination of LFSR and NFSR such as Grain-128a, it is possible to mount a divide-and-conquer attack. For example the attack in [54] on Grain-128a finds the state of the LFSR independently of the state of the NFSR. (In some ways this is reminiscent of the attack by Siegenthaler [53] on the nonlinear combiner model.) So even though Grain-128a uses a 256-bit state, due to the divide-and-conquer strategy the full protection of the large state is not achieved. On the other hand, for the nonlinear filter model, there is no known way to mount a divide-and-conquer attack. Our proposal at the 128-bit security level uses a 257-bit LFSR, and there is no way to estimate half the state of the LFSR without involving the other half.
@@ 2 Preliminaries
For $n\geq 0$, let $\mathbf{x}=(x_{1},\ldots,x_{n})$ be an $n$-bit binary string. The weight of $\mathbf{x}$ is ${\sf wt}(\mathbf{x})=\#\{i:x_{i}=1\}$. By $\mathbf{0}_{n}$ and $\mathbf{1}_{n}$ we will denote the all-zero and all-one strings of length $n$ respectively. Let $\mathbf{x}=(x_{1},\ldots,x_{n})$ and $\mathbf{y}=(y_{1},\ldots,y_{n})$ be two $n$-bit strings. The distance between $\mathbf{x}$ and $\mathbf{y}$ is $d(\mathbf{x},\mathbf{y})=\#\{i:x_{i}\neq y_{i}\}$; the inner product of $\mathbf{x}$ and $\mathbf{y}$ is $\langle\mathbf{x},\mathbf{y}\rangle=x_{1}y_{1}\oplus\cdots\oplus x_{n}y_{n}$; and we define $\mathbf{x}\leq\mathbf{y}$ if $x_{i}\leq y_{i}$ for $i=1,\ldots,n$.
@@
Let $f$ be an $n$-variable function. The algebraic normal form (ANF) representation of $f$ is the following: $f(X_{1},\ldots,X_{n})=\bigoplus_{\bm{\alpha}\in\mathbb{F}_{2}^{n}}a_{\bm{\alpha}}\mathbf{X}^{\bm{\alpha}}$, where $\mathbf{X}=(X_{1},\ldots,X_{n})$; for $\bm{\alpha}=(\alpha_{1},\ldots,\alpha_{n})\in\mathbb{F}_{2}^{n}$, $\mathbf{X}^{\bm{\alpha}}$ denotes $X_{1}^{\alpha_{1}}\cdots X_{n}^{\alpha_{n}}$; and $a_{\bm{\alpha}}\in\mathbb{F}_{2}$. The (algebraic) degree of $f$ is ${\sf deg}(f)=\max\{{\sf wt}(\bm{\alpha}):a_{\bm{\alpha}}=1\}$. Functions of degree at most 1 are said to be affine functions. Affine functions with $a_{\mathbf{0}_{n}}=0$ are said to be linear functions. It is known that if $f$ is balanced, then ${\sf deg}(f)\leq n-1$.
@@
For $n\geq 1$, let ${\sf Maj}:\{0,1\}^{n}\rightarrow\{0,1\}$ be the majority function defined in the following manner. For $\mathbf{x}\in\{0,1\}^{n}$, ${\sf Maj}(\mathbf{x})=1$ if and only if ${\sf wt}(\mathbf{x})\geq\lceil n/2\rceil$. It is known [24] that ${\sf Maj}_{n}$ has the maximum possible AI of $\lceil n/2\rceil$.

### v5

@@ Abstract
The nonlinear filter model is an old and well understood approach to the design of stream ciphers. Extensive research over several decades has shown how to attack stream ciphers based on this model and has identified the security properties required of the Boolean function used as the filtering function to resist such attacks. This led to the problem of constructing Boolean functions which provide adequate security and at the same time are efficient to implement. Unfortunately, over the last two decades no good solutions to this problem appeared in the literature. The lack of good solutions has effectively led to nonlinear filter model becoming more or less obsolete. This is a big loss to the cryptographic design toolkit, since the great advantage of the nonlinear filter model lies, beyond its simplicity and its ability to provide low-cost solutions for hardware-oriented stream ciphers, in the accumulated knowledge about the security requirements for the filtering function, which gives confidence in its security when all criteria are met. In this paper we construct balanced functions on an odd number $n\geq 5$ of variables with the following provable properties: linear bias equal to $2^{-\lfloor n/2\rfloor-1}$, algebraic degree equal to $2^{\lfloor\log_{2}\lfloor n/2\rfloor\rfloor}$, algebraic immunity at least $\lceil(n-1)/4\rceil$, fast algebraic immunity at least $1+\lceil(n-1)/4\rceil$, and the functions can be implemented using $O(n)$ NAND gates. The functions are obtained from a simple modification of the well known class of Maiorana-McFarland bent functions. Due to the efficiency of implementation, for any target security level, we can construct functions which provide the required level of resistance to fast algebraic and fast correlation attacks. By appropriately choosing $n$ and the length $L$ of the linear feedback shift register, we show that it is possible to obtain examples of stream ciphers which are provably $\kappa$-bit secure against well known classes of attacks for various values of $\kappa$. We provide concrete proposals for $\kappa=80,128,160,192,224$ and $256$ using LFSRs of lengths 163, 257, 331, 389, 449, 521 and filtering functions on 75, 119, 143, 175, 203 and 231 variables. For the $80$-bit, $128$-bit, and the $256$-bit security levels, the circuits for the corresponding stream ciphers require about 1743.5, 2771.5, and 5607.5 NAND gates respectively. For the $80$-bit and the $128$-bit security levels, the gate count estimates compare quite well to the famous ciphers Trivium and Grain-128a respectively, while for the $256$-bit security level, we do not know of any other stream cipher design which has such a low gate count.
@@ 1 Introduction
The nonlinear filter model of stream ciphers is several decades old; one may note that the model was extensively discussed in the book by Rueppel [49] which was published in the mid-1980s. The nonlinear filter model consists of two components, namely a linear feedback shift register (LFSR) and a Boolean function $f$ which is applied to a subset of the bits of the LFSR at fixed positions (called tap positions). At each clock, $f$ is applied to the present state of the LFSR to produce a single keystream bit and simultaneously the LFSR also moves to the next state. LFSRs are efficient to implement in hardware. So the implementation efficiency of the nonlinear filter model is essentially determined by the efficiency of implementation of $f$.
@@
A class of guess-then-determine attacks are known against the nonlinear filter model. The inversion attack [27] is the first such attack. Subsequently, a number of papers have developed the idea into the generalised inversion attack, the filter state guessing attack, and the generalised filter state guessing attack [27, 28, 33, 47, 58]. Recommendations for resistance to the state guessing attack do not specify conditions on the filtering function. Rather, the recommendations specify conditions on the tap positions, i.e. the LFSR positions where the filtering function inputs are drawn. We note that strictly following these recommendations with a LFSR of usual size (say about 256 flip-flops) implies taking a number of variables from the filtering function much lower than the desired security level.
@@
The well known Maiorana-McFarland class of bent functions is defined as follows. For $m\geq 1$, let $\mathbf{X}$ and $\mathbf{Y}$ be two vectors of $m$ variables. Then a $2m$-variable Maiorana-McFarland bent function is defined to map $(\mathbf{X},\mathbf{Y})$ to $\langle\pi(\mathbf{X}),\mathbf{Y}\rangle\oplus h(\mathbf{X})$, where $\pi$ is a bijection from $m$-bit strings to $m$-bit strings and $h$ is any $m$-variable Boolean function. It is well-known that the nonlinearity of Maiorana-McFarland functions does not depend on the choice of $h$ nor on that of the permutation $\pi$. Bad choices of both $\pi$ and $h$ (for example, $\pi$ to be the identity permutation and $h$ to be a constant function) provide functions whose algebraic immunity is low. On the bright side, we observe that it may be possible to improve the algebraic resistance by properly choosing $h$. Assume that $\pi$ is chosen to be the bit reversal permutation. It is known [24] that the majority function possesses the best possible algebraic immunity. Motivated by this fact, we choose $h$ to be the majority function on $m$ variables. We prove that the resulting bent function on $2m$ variables has algebraic immunity at least $\lceil m/2\rceil$, and hence fast algebraic immunity at least $1+\lceil m/2\rceil$. On the implementation aspect, we show that the majority function can be computed using $O(m)$ NAND gates. So the obtained bent function has maximum nonlinearity, sufficient (fast) algebraic immunity when the number of variables is large enough (to protect against fast algebraic attacks at a specified security level), and at the same time is quite efficient to implement. Indeed, the (fast) algebraic immunity is not the maximum possible, but due to the implementation efficiency, it is possible to increase the number of variables to achieve the desired level of algebraic resistance. A novelty of our work is the observation and the proof that choosing $h$ to be the majority function improves the algebraic immunity. While Maiorana-McFarland bent functions have been extensively studied in the literature, this simple observation has escaped the notice of earlier researchers.
@@
One problem is that a bent function is not balanced. This problem is easily rectified by XORing a new variable to the bent function, obtaining a function on an odd number of variables. This modification requires only one extra XOR gate for implementation. In terms of security, the modification does not change the linear bias. We prove that the algebraic immunity of the modified function is at least the algebraic immunity of the bent function, and hence the fast algebraic immunity is at least one plus the algebraic immunity of the bent function. So the algebraic resistance of the modified function is essentially the same as that of the bent function. A consequence of unbalancedness of the filtering function is that the keystream sequence is unbalanced, which leads to a distinguishing attack. The simple modification of XORing an extra variable prevents such a distinguishing attack. The resistance to fast correlation attacks and fast algebraic attacks are ensured by the bent function (on an appropriate number of variables). A positive aspect of XORing a new variable is that it prevents (see Theorem 2 of [27]) certain kinds of information leakage which was identified in [3]. On the negative side, filtering functions of the form $W+g(\mathbf{Z})$ were shown to be susceptible to the inversion attack [27]. We prevent this attack by choosing the gap between the first and the last tap positions to be sufficiently large; this countermeasure was already proposed in [27]. More generally, we show that due to the fact that we use filtering functions on a large number of variables, the various kinds of state guessing attacks [28, 33, 47, 58] do not apply to our proposals.
The literature [27, 28, 33, 47, 58] provides guidance on the choice of tap positions, i.e. the positions of the LFSR which are to be tapped to provide input to the filtering function. One such recommendation is to ensure that the tap positions form a full positive difference set. As we show later, this condition requires the length $L$ of the LFSR to be at least quadratic in the number $n$ of variables of the filtering function. In our proposals, the value of $n$ is close to the target security level $\kappa$. So if the tap positions are to be chosen to satisfy the full positive difference set condition, then $L$ will become too large, and the resulting stream cipher will be of no practical interest. The recommendations in the literature on selection of tap positions are essentially sufficient (but not necessary) conditions to prevent state guessing attacks. For our proposals, we show that such attacks can be avoided even though we do not follow the recommendations on tap positions.
We propose to use filtering function on a large number of variables. On the other hand, due to efficiency considerations, it is not possible to make the length of the LFSR too long. In particular, in all our concrete proposals, the ratio $n/L$ is close to half and is much higher than what was used in all previous proposals. Using a high value of $n/L$ means that the tap positions are placed much closer together than in previous proposals. This creates possibility of an overlap in the two sets of state bits determining two nearby keystream bits. Such overlap has the potential to cancel out terms and reduce the overall “complexity”. Our choice of tap positions is motivated by the requirement of ensuring that such cancellations do not take place. The generation of a keystream bit requires a call to the majority function. We set one of the goals of choosing the tap positions to be to ensure that the inputs to the two calls to majority for generating two different keystream bits have only a small overlap. In particular, we prove that for our choice of tap positions, there is no cancellation of terms of the majority function corresponding to two keystream bits. We evolved this design criterion for the tap positions in response to an attack [6] on an earlier version, and based on informal attack ideas (though not actual attacks) suggested by Subhadeep Banik, Willi Meier, and Bin Zhang also on a previous version. Further, we choose all the tap positions for the variables in $\mathbf{X}$ to the left of all the tap positions for the variables in $\mathbf{Y}$. This choice combined with the choice of $\pi$ as the bit reversal permutation allows us to prove that the quadratic terms arising from the application of the filtering function to the state bits corresponding to two different states do not cancel out with each other.
@@
We note that there are some old (prior to the advent of algebraic attacks) works [50, 51, 30] on efficient implementation of Boolean functions on a large number of variables targeted towards the nonlinear combiner model of stream ciphers. There are also a few later works [46, 18] on implementation of Maiorana-McFarland type functions. These works, however, do not cover the implementation of the functions that we introduce, the reason being that these new functions themselves do not appear earlier in the literature.
@@
Well known stream ciphers such as Trivium [9], Mickey [4], Grain-128a [1], SNOW [26] and Sosemanuk [5], use novel and ingenious ideas. Nonetheless, these are standalone designs. The nonlinear filter model, on the other hand, is a model for stream cipher design. Due to the simplicity of the nonlinear filter model, provable properties of the filtering function can be translated into provable resistance of the stream cipher against well known classes of attacks. In particular, for the proposals that we put forward, the provable linear bias of the filtering function translates to provable protection against a large class of fast correlation attacks, and the provable lower bound on the (fast) algebraic immunity translates to provable resistance against algebraic attacks. Stream ciphers based on either nonlinear finite state machines, or nonlinear feedback shift registers do not enjoy this advantage, i.e. for such stream ciphers it is very hard to obtain provable guarantees against various well known types of attacks. As an example, our proposal at the 128-bit security level generates keystream for which the best linear approximation has provable linear bias of $2^{-60}$, while at the same security level, for Sosemanuk [5] the best known [39] linear approximation has correlation $2^{-20.84}$, and it is not known whether there are approximations with higher correlations. Similarly, the time complexities of fast algebraic attacks against Trivium, Mickey, Grain-128a, SNOW and Sosemanuk are not known. It is believed that these stream ciphers can withstand fast algebraic attacks. In contrast, we are able to prove that at the appropriate security levels, the fast algebraic attack is ineffective against the new nonlinear filter model based stream cipher proposals that we put forward.
One advantage of using LFSRs is that it is possible to provably ensure that the LFSR has a maximum period. For stream ciphers based on NFSRs, such as Trivium and Mickey, such provable assurance is not available. For stream ciphers which use a combination of LFSR and NFSR such as Grain-128a, it is possible to mount a divide-and-conquer attack. For example the attack in [54] on Grain-128a finds the state of the LFSR independently of the state of the NFSR. (In some ways this is reminiscent of the attack by Siegenthaler [53] on the nonlinear combiner model.) So even though Grain-128a uses a 256-bit state, due to the divide-and-conquer strategy the full protection of the large state is not achieved. On the other hand, for the nonlinear filter model, there is no known way to mount a divide-and-conquer attack. Our proposal at the 128-bit security level uses a 257-bit LFSR, and there is knwon no way to estimate half the state of the LFSR without involving the other half.
@@
Important note.
An earlier version of the proposal was attacked [6] using a differential attack. In response, we have modified the proposal to resist the attack in [6]. Though the modified proposal described in this version of the paper resists the attack in [6], we do not have any proof that the present proposal resists all kinds of differential attacks. We note that while provable properties of the filtering function can be translated to provable resistance against certain classes of attacks, it is by no means true that these properties provide resistance against all classes of attacks. We welcome further analysis and investigation of other avenues of attacks on the concrete stream cipher proposals that we put forward in this paper.
@@ 2 Preliminaries
For $n\geq 0$, let $\mathbf{x}=(x_{1},\ldots,x_{n})$ be an $n$-bit binary string. The support of $\mathbf{x}$ is ${\sf supp}(\mathbf{x})=\{1\leq i\leq n:x_{i}=1\}$, and the weight of $\mathbf{x}$ is ${\sf wt}(\mathbf{x})=\#{\sf supp}({\sf supp}(\mathbf{x}))$. By $\mathbf{0}_{n}$ and $\mathbf{1}_{n}$ we will denote the all-zero and all-one strings of length $n$ respectively. Let $\mathbf{x}=(x_{1},\ldots,x_{n})$ and $\mathbf{y}=(y_{1},\ldots,y_{n})$ be two $n$-bit strings. The distance between $\mathbf{x}$ and $\mathbf{y}$ is $d(\mathbf{x},\mathbf{y})=\#\{i:x_{i}\neq y_{i}\}$; the inner product of $\mathbf{x}$ and $\mathbf{y}$ is $\langle\mathbf{x},\mathbf{y}\rangle=x_{1}y_{1}\oplus\cdots\oplus x_{n}y_{n}$; and we define $\mathbf{x}\leq\mathbf{y}$ if $x_{i}\leq y_{i}$ for $i=1,\ldots,n$.
@@
Let $f$ be an $n$-variable function. The algebraic normal form (ANF) representation of $f$ is the following: $f(X_{1},\ldots,X_{n})=\bigoplus_{\bm{\alpha}\in\mathbb{F}_{2}^{n}}a_{\bm{\alpha}}\mathbf{X}^{\bm{\alpha}}$, where $\mathbf{X}=(X_{1},\ldots,X_{n})$; for $\bm{\alpha}=(\alpha_{1},\ldots,\alpha_{n})\in\mathbb{F}_{2}^{n}$, $\mathbf{X}^{\bm{\alpha}}$ denotes the monomial $X_{1}^{\alpha_{1}}\cdots X_{n}^{\alpha_{n}}$; and $a_{\bm{\alpha}}\in\mathbb{F}_{2}$. The (algebraic) degree of $f$ is ${\sf deg}(f)=\max\{{\sf wt}(\bm{\alpha}):a_{\bm{\alpha}}=1\}$. Functions of degree at most 1 are said to be affine functions. Affine functions with $a_{\mathbf{0}_{n}}=0$ are said to be linear functions. It is known that if $f$ is balanced, then ${\sf deg}(f)\leq n-1$.
@@
For $n\geq 1$, let ${\sf Maj}_{n}:\{0,1\}^{n}\rightarrow\{0,1\}$ be the majority function defined in the following manner. For $\mathbf{x}\in\{0,1\}^{n}$, ${\sf Maj}(\mathbf{x})=1$ if and only if ${\sf wt}(\mathbf{x})>\lfloor n/2\rfloor$. Clearly ${\sf Maj}_{n}$ is a symmetric function. The following results were proved in [24].
Theorem 1 (Theorems 1 and 2 of [24]).
Let $n$ be a positive integer.
  1. 1.
    ${\sf Maj}_{n}$ has the maximum possible AI of $\lceil n/2\rceil$.
  2. 2.
    The degree of ${\sf Maj}_{n}$ is equal to $2^{\lfloor\log_{2}n\rfloor}$.
  3. 3.
    Any monomial occurring in the ANF of ${\sf Maj}_{n}$ has degree more than $\lfloor n/2\rfloor$.